OpenStudy (anonymous):

Given that f(x) = 5x − 10 and g(x) = x + 3, solve for f(g(x)) when x = −1. −30 0 10 12

jimthompson5910 (jim_thompson5910):

What is g(-1) equal to?

OpenStudy (anonymous):

idk im confused

jimthompson5910 (jim_thompson5910):

g(x) = x + 3 g(-1) = -1 + 3 ... replace EVERY x with -1 g(-1) = ???

OpenStudy (anonymous):

ohhh would it be 0?

jimthompson5910 (jim_thompson5910):

the final answer, yes hopefully you see that once you get g(-1) = 2, you plug that into f(x)

OpenStudy (anonymous):

yeah i get it now thanks alot really appricate

jimthompson5910 (jim_thompson5910):

you're welcome